# Integrating Agent-Based and Compartmental Models for Infectious Disease Modeling: A Novel Hybrid Approach

### v1

@@ 3.3.1 Population flow between model regions
where $n_{out}$ is the number of agents leaving M2. This ensures that the flow between the two regions remains symmetric. The lag of one time step prevents scenarios where more agents leave M2 than are present, thus avoiding population inconsistencies. Since M2 does not simulate spatial movement the $y$-coordinate at which an agent leaves M2 is unknown. For variant A, this coordinate is sampled from a uniform distribution. For variant B, the marginal of the density function (see Equation 1) is used to derive sampling weights.
@@ 3.3.2 Infection dynamics within the ODE area
$\Phi_{1}(t+1)$ $=I(t)+\frac{\beta\cdot S(t)\cdot I(t)}{N(t)}-\gamma\cdot I(t)dt$
@@ 3.3.3 Parameter selection
To effectively compare and couple M1 and M2, both models need to produce similar epidemic outcomes given identical starting conditions. Following the procedure and disease characteristics in Breitwieser et al. [29], we first determine the parameters $\beta$ and $\gamma$ for M2, corresponding to a measles epidemic with $R_{0}=12.9$ and an 8-day recovery period, resulting in $\beta=1.61$ and $\gamma=1/8$. Next, we generate a time series of infectious individuals by numerically solving the compartmental model using initial values of $S=995,I=5,R=0$ and a step size of 0.33 days. This is achieved with the RK45 method, as implemented in SciPy’s integrate.solve_ivp function [35].
We then fit the parameters $\Lambda$, $\rho$, and $\theta$ for M1-A to match the time series data, with the same initial conditions and a recovery rate of $1/24$ (to correspond to a step size of 0.33 days). This optimization is done using the Nelder-Mead minimizer (as implemented in SciPy’s optimize.minimize [35]) with least squares as the cost function, yielding $\Lambda=0.215$, $\rho=0.025$, and $\theta=1.162$.
We repeat the procedure for M1-B, and additionally fit the parameter $\delta$ (the base step size), fixing $\eta$ at 3, and positioning the attraction point at $X=Y=4.5$. This results in $\Lambda=0.268$, $\rho=0.016$, $\theta=1.01$, and $\delta=0.01$.
Figure 4 shows that M1-A and M2 produce closely aligned SIR curves, while M1-B exhibits a more rapid rise and a lower peak of infections. This is consistent with the finding in Susandi et al. [36] that ABMs with heterogeneous populations cannot produce identical results to compartmental models - which explains this discrepancy.
@@
In the following section, we describe the numerical experiments conducted to evaluate the performance and robustness of our hybrid modeling approach. The primary objective is to understand how varying the weight of the compartmental (ODE-based) and agent-based component affects model predictions, infection dynamics, and computational efficiency. By systematically adjusting these weights, we aim to uncover the nuanced behavior of the hybrid model under different configurations. This analysis will provide valuable insights into the potential applications and limitations of the suggested hybrid model in infectious disease modeling.
@@ 4 Experimental Setup and Methodology
In this section, we describe the experimental setup and methodology used to evaluate the fundamental dynamics of our hybrid model. By varying the weight between its two main components - the agent-based model (M1) and the compartmental model (M2) - we aim to understand how altering the balance between these components and the location of the interface affects the model outcomes in terms of the infectious wave. This setup allows us to test the robustness of our model under various conditions. The results of these experiments are presented in the subsequent section.
@@ 4.1 Model Weights
We adjust the weights of M1 ($w_{1}$) and M2 ($w_{2}$) by shifting the interface in the hybrid model along the x-axis. For example, given $0\leq x\leq 9$, placing the interface at x=4.5 results in $w_{1}=w_{2}=0.5$, while a placement at x=6.75 results in $w_{1}=0.25$ and $w_{2}=0.75$. The interface shift allows us to systematically study the impact of different model configurations on infection dynamics.
@@ 4.2 Model Configurations
We compare two configurations of the hybrid model: Hybrid Model A (M1-A with M2) and Hybrid Model B (M1-B with M2). This comparison allows us to examine our hybrid model under two scenarios. In the first scenario (Hybrid Model A), both model components simulate a homogeneous population, leading to very similar results individually. In the second scenario (Hybrid Model B), the agent-based component simulates a heterogeneous population, yielding different epidemic dynamics compared to the M2-component. In this scenario, the attraction point is placed at the center of the simulation space $(x=4.5,y=4.5)$, making the distance of the attraction point to the interface dependent on $w$. For $w_{2}>0.5$, the attraction point is located within the borders of M2. Additionally, we examine a third scenario in the appendix, where the location of the attraction point is always placed at the center of M1, rather than at the center of the full simulation space. These scenarios enable us to study the implications of spatial differences on spatial coupling effects.
@@ 4.3 Evaluation Metrics
[…]
@@ 5 Results
This section presents the results of our experiments, evaluating the performance and robustness of the hybrid model under different configurations. Specifically, we examine how varying the weight of the compartmental model (M2) and the agent-based model (M1) affects various metrics (see previous section) related to infection dynamics and computational efficiency in both Hybrid Model A (M1-A with M2) and Hybrid Model B (M1-B with M2).
Results are shown in Figures 5(a)-5(f) for Hybrid Model A and in Figures 6(a)-6(f) for Hybrid Model B. The x-axis represents the weight of M2 ($w_{2}$). When $w_{2}=0$, the entire population is modeled by M1, and when $w_{2}=1$, the entire population is modeled by M2.
Figure 5: Hybrid model A: Average results for 100 independent realizations per level of model 2 weight ($w_{2}$). Bars around points represent the standard deviation. (a) RMSE for the time series of infectious between the reference model and the hybrid model. (b) Peak time of the time series of infectious. (c) Peak height of the time series of infectious. (d) Fisher-Pearson coefficient of skewness of the time series of infectious. (e) FWHM of the time series of infectious. (f) CPU time in seconds.
@@ 5.1.2 Peak height
The peak height of the infection curve does not change significantly with changing $w_{2}$. It remains around 71.57% at $w_{2}=0$ and 72.46% at $w_{2}=1$. This consistency is expected since the peak height is similar in both individual models M1-A and M2, as shown in 4. Similar to the RMSE, the standard deviation shrinks with increasing $w_{2}$.
@@ 5.1.3 Peak time
A notable coupling effect is observed in the peak time of the infectious curve. While the peak time is slightly earlier in the reference model ($16t$) compared to M1-A ($17.36t$), it does not gradually decrease as $w_{2}$ increases. Instead, there seems to be a quadratic relationship between the peak time and $w_{2}$. This may be due the discrete implementation of time in our models and the relatively small difference in peak time of 1.36 time steps between the model components, causing the resulting peak time to align more closely with the model component modeling the majority of the population.
@@ 5.1.6 CPU time
The CPU time decreases from 3.18 to 0.09 seconds as $w_{2}$ increases, following a convex pattern. This diminishing reduction in CPU time can be attributed to the stochastic processes computed for each agent, where calculations often involve interactions with all other agents. Thus, the reduction in computational load is more substantial when increasing $w_{2}$ from lower values (e.g., 0 to 0.02) compared to higher values (e.g., 0.9 to 0.92), reflecting the non-linear scaling of interaction complexity.
@@ 5.2 Hybrid model B
Figure 6: Hybrid model B: Average results for 100 independent realizations per level of model 2 weight ($w_{2}$). Bars around points represent the standard deviation. (a) RMSE for the time series of infectious between the reference model and the hybrid model. (b) Peak time of the time series of infectious. (c) Peak height of the time series of infectious. (d) Fisher-Pearson coefficient of skewness of the time series of infectious. (e) FWHM of the time series of infectious. (f) CPU time in seconds.
@@ 5.2.1 Root Mean Squared Error (RMSE)
The RMSE at $w_{2}=0$ is 44.3, nearly twice as large as the RMSE for M1-A. Similar to Hybrid Model A, the RMSE for Hybrid Model B drops towards zero as $w_{2}$ increases. However, there is a notable minimum around the 0.5-mark of $w_{2}$. The RMSE decreases for $w_{2}$ values below 0.5, then increases for $w_{2}$ values above 0.5 before dropping again after the 0.6-mark. This indicates that M2 has a stronger effect on model results around the 0.5 mark than at the 0.6 mark. This can be attributed to the attraction point in the agent-based component, resulting in a spatially heterogeneous population. Agents form a cluster where the infection spreads more rapidly. When $w_{2}$ is set to 0.5, the attraction point lies on the interface between M1-B and M2, attracting many agents into M2 and therefore increasing the interaction between the models.
Figure 7: Changes in population metrics for Hybrid Model A and B for different weights of M2. Results from 100 independent realizations. (a) Average percentage of population entering M2 per $t$. (b) Average percentage of overall population in M1 per $t$.
Indeed, Figure 7(a) shows that the average population flow from M1 to M2 during a time step depends on $w_{2}$. In Hybrid Model A, it is around 1.5% of the total population for weights below 0.8, before dropping for higher weights. In Hybrid Model B, it increases from 0.02 to 0.53, peaking at 4.5%, before decreasing rapidly for higher weights. Additionally, Figure 7(b) shows that the share of the overall population present in M1 drops faster around $w=0.5$ in Hybrid Model B, while decreasing steadily in Hybrid Model A.
@@ 5.2.3 Peak time
The peak time of a full M1-B is $16.66t$ (at $w_{2}=0$), slightly closer to the reference model than M1-A. The attraction point appears to influence the peak time of the hybrid model. In the full M1-B model, the peak is delayed by 0.66 compared to M2. In Hybrid Model B, the peak time is closer to the full M2 result around $w_{2}=0.4$ than around $w_{2}=0.6$, again likely due to the increased between-model interaction around the 0.5-mark. The local minimum around the 0.4 mark, rather than the 0.5 mark, suggests that the discrete nature of time amplifies the interaction effect between the model components.
@@ 5.2.5 Full-Width-at-Half-Maximum (FWHM)
At $w_{2}=0$, the FWHM is considerably higher (29.8) than at $w_{2}=1$ (25.0), indicating that the infectious peak of M1-B is more truncated than in M1-A or M2. Similar to the previous metric, a regression can be observed as $w_{2}$ increases with a temporary shift after the 0.38-mark. Both observations are consistent with the increased between-model exchange around this value.
@@ 5.2.6 CPU time
For lower $w_{2}$-values, the average CPU time of Hybrid Model B is higher than in the case of Hybrid Model A, at 3.37 seconds when $w_{2}=0$. This is due to the increased complexity introduced by the attraction points. Additionally, there is an increase from $w_{2}=0$ to $w_{2}=0.06$, possibly due to the further complexity introduced by population flow from M2 to M1, influenced by the landscape in M1. For Hybrid Model B, the drop in CPU time follows the drop in population size of the M1-B component, as visualized in Figure 7(b). The computational demands of M1 stem from simulating each individual agent, making the population size a key factor in determining CPU time.
@@ 6 Discussion
In contrast, the results for Hybrid Model B reveal a more complex interaction involving the implemented attraction point and the between-model interface. This interaction triggers switches in the model outcomes, favoring one component over the other under different conditions. The direction-switching effect observed in the peak time and peak height for Hybrid Model B can be explained by the presence of the attraction point. Without the attraction point, these metrics would likely follow a trajectory similar to Hybrid Model A. However, the increased between-model interaction around $w_{2}=0.5$ leads to a local change, emphasizing the sensitivity of the hybrid model to the spatial configuration of the attraction point.
@@ 6.1 Impact on Infection Spread Dynamics
The study reveals that spatial coupling of an ABM with an ODE-based model can effectively capture the dynamics of infection spread in the overall population. In Hybrid Model A, where both components simulate a homogeneous population, the infection dynamics transition smoothly as the weight shifts between M1 and M2. This smooth transition reflects a successful integration of the detailed agent interactions in M1 with the aggregate population-level processes in M2, maintaining consistency in the model’s outcomes.
@@
These findings highlight that the location of the coupling mechanism in a hybrid model can amplify the differences between model components. Differences between models and sub-populations, as well as the level of population flow, should be carefully considered in real-world applications. The increased between-model interaction observed in Hybrid Model B underscores the need for a nuanced approach to spatial coupling, especially when dealing with heterogeneous populations, thereby enhancing the model’s applicability and precision in various scenarios.

### v2

@@ 3.3.1 Population Flow between Model Regions
where $n_{out}$ is the number of agents leaving M2. This ensures that the flow between the two regions remains symmetric. The lag of one time step prevents scenarios where more agents leave M2 than are present, thus avoiding population inconsistencies. Since M2 does not simulate spatial movement, the $y$-coordinate at which an agent leaves M2 is unknown. For M1-A, this coordinate is sampled from a uniform distribution. For M1-B, the marginal of the density function (see Equation 1) is used to derive sampling weights.
@@ 3.3.2 Infection Dynamics within the ODE Area
$\Phi_{1}(t+1)$ $=I(t)+\frac{\beta\cdot S(t)\cdot I(t)}{N(t)}dt-\gamma\cdot I(t)dt$
@@ 3.3.3 Parameter Selection
To effectively compare and couple M1 and M2, both models need to produce similar epidemic outcomes given the same starting conditions. However, an ABM featuring a heterogeneous population, as is the case with M1-B, cannot produce identical results to an analogous compartmental model [39]. Our aim with M1-B is therefore to achieve the closest feasible alignment with the M2 outcomes.
Producing comparable dynamics can be achieved by parametrizing one model first and adjusting the other model to match. Rahmandad and Sterman [40] argue that in the real world, where contact structures are often unknown, compartmental model parameters are usually fitted to epidemic data, effectively capturing the impact of network topology and individual variability. However, with known values for the basic reproduction number ($R_{0}$) and disease duration, parametrizing a compartmental model first and fitting the ABM parameters afterwards simplifies calibration and requires fewer assumptions. We thus adopted this latter approach and followed the methodology and disease characteristics for a measles epidemic outlined in Breitwieser et al. [33], allowing us to focus on evaluating the hybrid coupling mechanism without introducing additional complexity.
Using parameters $\beta=1.61$ and $\gamma=1/8$ for a measles epidemic with $R_{0}=12.9$ and an 8-day recovery period, we generated a time series of infectious individuals with M2. This was done by numerically solving the compartmental model with initial values of $S=995,I=5,R=0$ and a step size of 0.33 days, using the RK45 method as implemented in SciPy’s integrate.solve_ivp function [41].
We then fitted the parameters $\Lambda$, $\rho$, and $\theta$ for M1-A to match the time series data, with the same initial conditions and a recovery rate of $1/24$ (corresponding to a step size of 0.33 days). This optimization was done using the Nelder-Mead minimizer (as implemented in SciPy’s optimize.minimize [41]) with least squares as the cost function, yielding $\Lambda=0.215$, $\rho=0.025$, and $\theta=1.162$.
We repeated the procedure for M1-B, and additionally fitted the parameter $\delta$ (the base step size), fixing $\eta$ at 3, and positioning the attraction point at $X=Y=4.5$. This results in $\Lambda=0.268$, $\rho=0.016$, $\theta=1.01$, and $\delta=0.01$.
Figure 4 shows that M1-A and M2 produce closely aligned SIR curves, while M1-B exhibits a more rapid rise and a lower peak of infections. This discrepancy is expected because, under population heterogeneity, infections quickly spread in dense clusters. As a result, the overall pool of susceptible individuals decreases faster, leading to an accelerated initial rise and a reduced overall dispersion of infections [40].
@@
In the following section, we describe the numerical experiments conducted to evaluate the performance and robustness of our hybrid modeling approach. The primary objective is to understand how varying the weight of the compartmental (ODE-based) and agent-based component affects model predictions, infection dynamics, and computational efficiency. By systematically adjusting these weights, we aim to uncover the nuanced behavior of the hybrid model under different configurations. This analysis provides valuable insights into the potential applications and limitations of the suggested hybrid model in infectious disease modeling.
@@ 4 Experimental Setup and Methodology
In this section, we describe the experimental setup and methodology used to evaluate the fundamental dynamics of our hybrid model. By varying the weight between its two main components - the ABM (M1) and the compartmental model (M2) - we aim to understand how altering the balance between these components and the location of the interface affects model outcomes in terms of the infectious wave. This setup allows us to test the robustness of our model under various conditions. The results of these experiments are presented in the subsequent section.
@@ 4.1 Model Weights
We adjusted the weights of M1 ($w_{1}$) and M2 ($w_{2}$) by shifting the interface in the hybrid model along the x-axis. For example, given $0\leq x\leq 9$, placing the interface at x=4.5 results in $w_{1}=w_{2}=0.5$, while a placement at x=6.75 results in $w_{1}=0.25$ and $w_{2}=0.75$. The interface shift allows us to systematically study the impact of different model configurations on infection dynamics.
@@ 4.2 Model Configurations
We compare two configurations of the hybrid model: Hybrid Model A (M1-A with M2) and Hybrid Model B (M1-B with M2), enabling us to examine its performance under two scenarios. In the first scenario, both model components of Hybrid Model A simulate a homogeneous population, leading to very similar results individually. In the second scenario, Hybrid Model B’s agent-based component simulates a heterogeneous population, yielding distinct epidemic dynamics compared to the M2-component. In this scenario, the attraction point is placed at the center of the simulation space $(x=4.5,y=4.5)$, making the distance of the attraction point to the interface dependent on $w$. For $w_{2}>0.5$, the attraction point falls within the borders of M2. Additionally, we examine a third scenario in the appendix, where the the attraction point is always placed at the center of M1, rather than at the center of the full simulation space. These scenarios enable us to study the implications of spatial differences for spatial coupling.
@@ 4.3 Evaluation Metrics
For the experiments, we evaluate the following metrics for the resulting time series of the infectious population: (a) the root mean squared error (RMSE) compared to the baseline from M2, quantifying the overall deviation from the reference model; (b) the peak height; (c) the time step at which the peak occurs; (d) the Fisher-Pearson coefficient of skewness, indicating the buildup speed of the infectious wave; (e) the full-width-at-half-maximum (FWHM), indicating the sharpness of the peak. Metrics b - e provide detailed insights into the dynamics and shape of the epidemic. Finally, we measure (f) the average CPU time for each model configuration. We increase the weight of M2 from 0 to 1 in 0.02 steps, averaging the metrics after 100 independent realizations per configuration.
@@ 5 Results
This section presents the results of our experiments, evaluating the performance and robustness of the hybrid model under different configurations. Specifically, we examine how varying the weight of the ABM (M1) and the compartmental model (M2) affects various metrics (see previous section) related to infection dynamics and computational efficiency in both Hybrid Model A (M1-A with M2) and Hybrid Model B (M1-B with M2). Results are shown in Figures 5(a)-5(f) for Hybrid Model A and in Figures 7(a)-7(f) for Hybrid Model B. The x-axis represents the weight of M2 ($w_{2}$). When $w_{2}=0$, the entire population is modeled by M1, and when $w_{2}=1$, the entire population is modeled by M2.
Figure 5: Hybrid Model A: Average results for 100 independent realizations per M2 weight ($w_{2}$). Bars around points represent the standard deviation. (a) RMSE for the infectious time series between the reference model and the hybrid model. (b) Peak time of the infectious time series. (c) Peak height of the infectious time series. (d) Fisher-Pearson coefficient of skewness of the infectious time series. (e) FWHM of the infectious time series. (f) CPU time in seconds.
@@ 5.1.2 Peak Height
The peak height of the infection curve does not change significantly across $w_{2}$-levels. It remains around 71.57% at $w_{2}=0$ and 72.46% at $w_{2}=1$. This consistency is expected since the peak height is similar in both individual models M1-A and M2, as shown in Figure 4. Similar to the RMSE, the standard deviation shrinks with increasing $w_{2}$.
@@ 5.1.3 Peak Time
A notable coupling effect is observed in the peak time of the infectious curve. While the peak occurs slightly earlier in the reference model ($16t$) compared to M1-A ($17.36t$), it does not gradually decrease as $w_{2}$ increases. Instead, there seems to be a quadratic relationship between the peak time and $w_{2}$. This may be due to the discrete implementation of time in our models and the relatively small difference in peak time of 1.36 time steps between the model components, causing the resulting peak time to align more closely with the model component modeling the majority of the population.
@@ 5.1.6 CPU Time
The CPU time decreases from 3.18 to 0.09 seconds as $w_{2}$ increases, following a convex pattern. This diminishing reduction in CPU time can be attributed to the stochastic processes computed for each agent, where calculations often involve interactions with all other agents. Thus, the reduction in computational load is more substantial when $w_{2}$ increases from lower values (e.g., 0 to 0.02) compared to higher values (e.g., 0.9 to 0.92), reflecting the non-linear scaling of interaction complexity.
@@ 5.2.1 Root Mean Squared Error (RMSE)
The RMSE at $w_{2}=0$ is 44.3, nearly twice as large as the RMSE for M1-A. Similar to Hybrid Model A, the RMSE for Hybrid Model B drops towards zero as $w_{2}$ increases. However, there is a notable minimum around the 0.5-mark of $w_{2}$. The RMSE decreases for $w_{2}$ values below 0.5, then increases for $w_{2}$ values above 0.5 before dropping again after the 0.6-mark. This indicates that M2 has a stronger effect on model results around the 0.5 mark than at the 0.6 mark. This can be attributed to the attraction point in the agent-based component, resulting in a spatially heterogeneous population: agents form a cluster where the infection spreads more rapidly. When $w_{2}$ is set to 0.5, the attraction point lies on the interface between M1-B and M2, attracting many agents into M2 and therefore increasing the interaction between the models.
Figure 6: Changes in population metrics for Hybrid Model A and B for different M2 weights ($w_{2}$). Results from 100 independent realizations. (a) Average percentage of population entering M2 per $t$. (b) Average percentage of total population in M1 per $t$.
Figure 7: Hybrid Model B: Average results for 100 independent realizations per M2 weight ($w_{2}$). Bars around points represent the standard deviation. (a) RMSE for the infectious time series between the reference model and the hybrid model. (b) Peak time of the infectious time series. (c) Peak height of the infectious time series. (d) Fisher-Pearson coefficient of skewness of the infectious time series. (e) FWHM of the infectious time series. (f) CPU time in seconds.
Indeed, Figure 6(a) shows that the average population flow from M1 to M2 during a time step depends on $w_{2}$. In Hybrid Model A, it is around 1.5% of the total population for weights below 0.8, before dropping for higher weights. In Hybrid Model B, it increases from 0.02 to 0.53, peaking at 4.5%, before decreasing rapidly for higher weights. Additionally, Figure 6(b) shows that the share of the total population present in M1 drops faster around $w=0.5$ in Hybrid Model B, while decreasing steadily in Hybrid Model A.
@@ 5.2.3 Peak Time
The peak time of a full M1-B ($w_{2}=0$) is $16.66t$, having a delay of 0.66 that is slightly closer to the reference model than M1-A’s peak. The attraction point appears to influence the peak time of the Hybrid Model B: The peak time is closer to the reference model around $w_{2}=0.4$ than around $w_{2}=0.6$, again likely due to the increased between-model interaction near the 0.5-mark. The local minimum near $w_{2}=0.4$ rather than around $w_{2}=0.5$ suggests that the discrete nature of time amplifies the interaction effect between the model components.
@@ 5.2.5 Full-Width-at-Half-Maximum (FWHM)
At $w_{2}=0$, the FWHM is considerably higher (29.8) than at $w_{2}=1$ (25.0), indicating that the peak in M1-B is more truncated than in M1-A or M2. Similar to the previous metric, a regression can be observed as $w_{2}$ increases with a temporary shift after the 0.38-mark. Both observations are consistent with the increased between-model exchange near this value.
@@ 5.2.6 CPU Time
For lower $w_{2}$-values, the average CPU time is higher for Hybrid Model B compared to Hybrid Model A, at 3.37 seconds when $w_{2}=0$. This is due to the increased complexity introduced by the attraction points. Additionally, there is an increase from $w_{2}=0$ to $w_{2}=0.06$, possibly due to the further complexity introduced by population flow from M2 to M1, influenced by the landscape in M1. For Hybrid Model B, the drop in CPU time follows the drop in population size of the M1-B component, as visualized in Figure 6(b). The computational demands of M1 stem from simulating each individual agent, making the population size a key factor in determining CPU time.
@@ 6 Discussion
In contrast, the results for Hybrid Model B reveal a more complex interaction involving the landscape-driven agent movement and the between-model interface. This interaction triggers switches in the model outcomes, favoring one component over the other under different conditions. The direction-switching effect observed in the peak time and peak height for Hybrid Model B can be explained by the presence of the attraction point. Without the attraction point, these metrics would likely follow a trajectory similar to Hybrid Model A. However, the increased between-model interaction around $w_{2}=0.5$ leads to a local change, emphasizing the sensitivity of the hybrid model to the spatial configuration of the attraction point.
@@ 6.1 Impact on Infection Spread Dynamics
The study reveals that spatial coupling of an ABM with an ODE-based model can effectively capture the dynamics of infection spread in the overall population. In Hybrid Model A, where both components simulate a homogeneous population, the infection dynamics transition smoothly as the weight shifts between M1 and M2. This smooth transition reflects a successful integration of the detailed agent interactions in M1 with the aggregate population-level processes in M2, maintaining consistency in model outcomes.
@@
These findings highlight that the location of the coupling mechanism in a hybrid model can amplify the differences between model components. Differences between sub-models and sub-populations, as well as the level of population flow, should be carefully considered in real-world applications. The increased between-model interaction observed in Hybrid Model B underscores the need for a nuanced approach to spatial coupling, especially when dealing with heterogeneous populations, thereby enhancing the model’s applicability and precision in various scenarios.